if i have 1200 feet of fencing what is the largest rectangular area i could have if i wanted to fence my backyard using the side of my house as a side?
Answer by venugopalramana(3286)

if i have 1200 feet of fencing what is the largest rectangular area i could have if i wanted to fence my backyard using the side of my house as a side
FENCING IS NEEDED FOR 3 SIDES ONLY
LENGTH OF FENCING = L+L+B=2L+B=1200
B=1200-2L
AREA=A=L*B=L(1200-2L)=1200L-2L^2=-2[L^2-600L)
A=-2[L^2-2(L)300+300^2-300^2]
=-2(L-300)^2+2*90000
A=180000-2(L-300)^2
A WILL BE MAXIMUM WHEN L=300
MAXIMUM FENCING AREA=180000